OpenStudy (anonymous):

If f(x) = 5x – 3, then what is f–1(x) ?

OpenStudy (across):

(x+3)/5

OpenStudy (anonymous):

f(-1)?

OpenStudy (anonymous):

or f(-x)?

OpenStudy (anonymous):

You mean the inverse of f(x). Just swap your x and y variables and solve for y.

OpenStudy (ash2326):

@billybob3 did you understand how to get this?

OpenStudy (anonymous):

i think he's just waiting for the answer :(

OpenStudy (anonymous):

yeah, you switch x and y

OpenStudy (anonymous):

Yes, so what's the problem?

OpenStudy (ash2326):

@billybob3 Inverse of a function is defined such that \[f^{-1} (f(x))=f(f^{-1}(x))=x\] We'll use this to find \(f^{-1} (x)\) We have \[f(x)=5x-3\] to find \(f^{-1} (x)\) put x= \(f^{-1} (x)\), we have \[f(f^{-1}(x))= 5 f^{-1} (x)-3\] we get \[x= 5 f^{-1} (x)-3\] \[5 f^{-1} (x)= x+3\] or \[f^{-1} (x)=\frac{x+3}{5}\]